OpenStudy (anonymous):

The velocity of a car traveling in a straight line increases from 0 meters/second to 30 meters/second in 8 seconds. What is the average acceleration of the car?

OpenStudy (xishem):

Acceleration is defined as: \[\left| \vec{a} \right|=\frac{\Delta v}{\Delta t}=\frac{v_f-v_i}{t_f-t_i}\]Does that help?